# Towards Mitigating Excessive Forgetting in LLM Unlearning via Entanglement-Aware Unlearning with Proxy Constraint

### v1

@@ 3.2 Loss Functions for Unlearning
    $L_{\text{GA}}(\theta^{\prime},\mathcal{D}_{f})=-\mathbb{E}_{\mathcal{D}_{f}}[\log(f(\theta^{\prime},y|x))],$ (1)
    where the optimization is performed on the forget dataset $\mathcal{D}_{f}$ without referencing the retain dataset $\mathcal{D}_{r}$. Notably, the target label $y$ used in this objective can be flexibly chosen. Instead of the original correct output, one may substitute it with an arbitrary response $\widetilde{y}$ that the model is encouraged to output post-unlearning, such as random answers [6], or simply answering "I don’t know" [18].
@@
    $L(\theta^{\prime},\mathcal{D}_{r})=-\mathbb{E}_{\mathcal{D}_{r}}[\log(f(\theta^{\prime},y|x))],$ (2)
@@
  • Negative Preference Optimization (NPO): NPO performs preference optimization using only negative responses as supervision signals. The forget loss is formulated as:
    $L_{\text{NPO},\beta}(\theta^{\prime})=-\frac{2}{\beta}\mathbb{E}_{\mathcal{D}_{f}}\left[\log\sigma(-\beta\log\frac{f(\theta^{\prime},y|x)}{f(\theta^{o},y|x)})\right],$ (3)
    where $\sigma(t)=1/(1+e^{-t})$ is the sigmoid function, $\beta>0$ is the inverse temperature and $f(\theta^{o})$ is the model trained on $\mathcal{D}_{\text{train}}$. NPO+GD combines the forgetting loss from NPO with the retaining loss from GD.
3.3 Memorization Score
Deep neural networks are known to “memory" their training data. In particular, Feldman et al. [8] show that instance-level memory is both common and essential under long-tailed data distributions. In such settings, retaining specific labels plays a crucial role in achieving near-optimal generalization. This perspective underpins the use of memorization scores below as a proxy to quantify the degree to which individual samples are preserved in the model’s memory.
Assumption 1 (Memorization Score[8]).
@@ 3.3 Memorization Score
$\text{mem}(\mathcal{A},\mathcal{D},i)=\Pr_{f\sim\mathcal{A}(\mathcal{D})}\left[f(x_{i})=y_{i}\right]-\Pr_{f\sim\mathcal{A}(\mathcal{D}\setminus{i})}\left[f(x_{i})=y_{i}\right],$ (4)
@@
4 Limitations of Existing Methods and Our Insights
Neural networks do not memorize all samples uniformly—some are consistently learned, while others are easily forgotten [9, 24]. This heterogeneity calls for sample-specific forgetting effort. To address this, Huang et al. [10] addressed this by assigning forgetting coefficients inversely proportional to the empirical loss of each forget sample.
Relying solely on empirical risk is not enough. The underlying intuition is that samples with lower loss have been more thoroughly memorized by the model and therefore should be more aggressively unlearned. However, this approach has a key limitation: empirical loss alone is an imperfect proxy for memorization depth. It merely reflects the model’s current predictive error on the forget data, but fails to capture how well the retrain model without the forget data would still generalize to that sample. As a result, relying solely on the current loss can misguide the unlearning process, since it fails to account for how the retain dataset may still enable the model to generalize to forget samples, making it unclear which examples truly require a stronger forgetting effort.
Memorization score enhances unlearning performance. Zhao et al. [35] introduce an ordered unlearning strategy guided by memorization score, which quantifies how well the model memorizes the data, providing a more faithful measure of memorization depth than empirical loss alone. By ranking forget samples according to these memorization scores, forget samples are unlearned sequentially with tailored unlearning methods and hyper-parameters, enabling differentiated forgetting effort across samples, effectively reducing collateral damage to retain knowledge and achieving a better trade-off between forgetting performance and model utility.
Figure 1: Overview of the EAGLE framework, which combines entanglement-awareness guided loss reweighting with Proxy Constraint to steer the unlearning process. This approach adaptively prioritizes forgetting of less entangled samples while leveraging proxy constraints to prevent over-forgetting, thereby improving Forget Quality and preserving model utility.
Memorization-based strategies come with high computational cost. However, despite its theoretical soundness and empirical benefits, this memorization-based strategy has a critical drawback: computing memorization scores requires access to the retrain model that excludes all forget data, which is often infeasible in LLMs. This contradicts the core motivation of fine-tuning-based unlearning, which is to avoid the substantial computational cost of retraining from scratch. This tension highlights the need for a scalable alternative that can approximate memorization without expensive retraining.
Memorization strength can be approximated by the entanglement between forget and retain samples. Furthermore, large language models (LLMs) inherently possess strong generalization capabilities, often achieving reasonable prediction accuracy on unseen inputs. This generalization arises from their ability to recognize and leverage semantic similarities between new inputs and previously learned data. As a result, certain forget samples may still be answered correctly, not because the model memorized them directly, but because they are entangled with retain dataset that shares latent patterns, phrasing, or structure. This retrieval-like behavior implies that forgetting such samples may require disproportionately more effort, as their representation is distributed across shared subspaces with retain data. Building on this insight, we argue that, instead of relying on costly retraining to measure memorization strength, one can approximate it by quantifying the degree of entanglement between each forget sample and the retain dataset in the embedding space.
5 Methodology
@@
Figure 2: Normalized distributions of different unlearning strategies and memorization score. All distributions are min-max normalized for fair comparison. EAGLE most closely approximates the memorization score distribution.
Figure 1 illustrates the overall workflow of EAGLE-PC: Entanglement-Aware Guided Loss REweighting with Proxy Constraint. Specifically, the entanglement-awareness guided loss reweighting adaptively modulates the forgetting effort for each example based on the similarity between forget and retain samples in the embedding space, prioritizing stronger forgetting effort for data that requires more unlearning (namely, less entangled samples). Meanwhile, the proxy data generated by GPT through in-context learning provides additional constraints that prevent over-forgetting of crucial knowledge, thereby enhancing the model utility and maintaining performance on retain information.
5.1 Entanglement-Awareness Guided Loss Reweighting (EAGLE)
In this paper, we propose a scalable method: Entanglement-Aware Guided Loss REweighting (EAGLE), a principled framework that enables fine-grained control over the forgetting process. Instead of relying on costly retraining to obtain memorization scores, we approximate them by measuring the cosine similarity between each forget sample’s embedding and the aggregated embedding of the retain dataset. Intuitively, forget samples that are more entangled with the retain knowledge—i.e., have higher embedding similarity are likely to be predicted more accurately by the retrain model. Thus, less forgetting effort is needed to match the performance of the retrain model. Conversely, samples that are less entangled can be forgotten more aggressively. By distributing the forgetting effort adaptively based on entanglement, we enable a fine-grained forgetting process that is both computationally efficient and aligned with the structure of learned knowledge.
Figure 2 compares the normalized L2 norms of forget-sample gradients from various approaches against the memorization score distribution, a widely recognized proxy for forgetting difficulty. Each distribution reflects the method’s weighting coefficient multiplied by the corresponding gradient norm, representing the sample’s effective contribution to the forgetting loss. Specifically, the vanilla method assigns a uniform weight to each forget sample, while EAGLE applies an entanglement-aware coefficient, as detailed in Algorithm 2. Another baseline, $1/loss^{0.5}$, follows the per-sample forgetting coefficient approach proposed by [10]. The distribution produced by EAGLE closely matches the memorization score distribution, while avoiding the high computational cost of directly calculating memorization scores. EAGLE reweights the loss contributions of forget samples by quantifying their entanglement with the retain set, measured via embedding similarity. Importantly, this approach does not require access to individual retain samples; it only relies on the average embedding of the retain dataset. If the model provider stores the overall training set embedding in advance, the retain embedding can be efficiently computed by subtracting the embedding of the forget set upon receiving a deletion request.
@@ 5.1 Entanglement-Awareness Guided Loss Reweighting (EAGLE)
Forget samples that are less entangled with retain dataset are typically semantically distant in the embedding space. These samples encode unique information with little overlap with the retain data, warranting stronger forgetting effort. EAGLE leverages this entanglement signal to guide unlearning, approximating it via the cosine similarity between each forget sample and the aggregated embedding of the retain dataset, and adaptively modulating forgetting effort for more precise and effective removal of target knowledge.
The overall framework is detailed in Algorithm 1, with detailed coefficient calculation in Algorithm 2. To address potential over-forgetting, we further introduce a penalty mechanism that leverages ICL-generated test data to constrain excessive forgetting, as detailed in Algorithm 3. Notably, the framework can be seamlessly integrated into existing unlearning methods, including gradient ascent [11], gradient difference [31], negative preference optimization (NPO) [34], and NPO+GD (NPO with gradient descent on the retain dataset).
Algorithm 1 EAGLE-PC unlearning framework
@@
1:Input: Retain dataset $\mathcal{D}_{r}=\left\{z_{j}\right\}_{j=1}^{|\mathcal{D}_{r}|}=\left\{(x_{j},y_{j})\right\}_{j=1}^{|\mathcal{D}_{r}|}$, Forget dataset $\mathcal{D}_{f}=\left\{z_{i}\right\}_{i=1}^{|\mathcal{D}_{f}|}=\left\{(x_{i},y_{i})\right\}_{i=1}^{|\mathcal{D}_{r}|}$, model loss function $f(\theta)$, retain strength $\alpha$,optimizer type opt,
@@
4:for $\left\{x_{i}.y_{i}\right\}\in\mathcal{D}_{r}$ do
5:  $E_{i}^{r}\leftarrow\text{AvgTokenEmbedding}(\theta_{0},x_{i})$
6:end for
[…]
8:for $t=0$ to T1T-1 do
9:  Sample $\xi_{t}=\left\{z_{i}\right\}_{i=1}^{m}$ from $\mathcal{D}_{f}$
10:  $w_{i}\leftarrow EAGLE(E_{\mathcal{D}_{r}},\xi_{t})$
11:  if opt == GA then
12:   $L(\theta_{t})\leftarrow-\frac{1}{m}\sum_{i=1}^{m}w_{i}\cdot f(\theta_{t},z_{i})+\mathcal{P}(\xi_{t},\mu)$
13:  else if opt == GD then
14:   Sample $\zeta_{t}=\{z_{j}\}_{j=1}^{n}$ from $\mathcal{D}_{r}$
15:   $L(\theta_{t})\leftarrow-\frac{1}{m}\sum_{i=1}^{m}w_{i}\cdot f(\theta_{t},z_{i})+\alpha\cdot\frac{1}{n}\sum_{j=1}^{n}f(\theta_{t},z_{j})+\mathcal{P}(\xi_{t},\mu)$
16:  else if opt == NPO then
17:   $\Delta_{i}\leftarrow w_{i}\cdot\left(f(\theta_{t},z_{i})-f(\theta_{0},z_{i})\right)$
18:   $L(\theta_{t})\leftarrow-\frac{2}{\beta}\cdot\frac{1}{m}\sum_{i=1}^{m}\log\sigma(\beta\cdot\Delta_{i})+\mathcal{P}(\xi_{t},\mu)$
19:  else if opt == NPO+GD then
20:   Sample $\zeta_{t}=\{z_{j}\}_{j=1}^{n}$ from $\mathcal{D}_{r}$
21:   $\Delta_{i}\leftarrow w_{i}\cdot\left(f(\theta_{t},z_{i})-f(\theta_{0},z_{i})\right)$
22:   $L(\theta_{t})\leftarrow-\frac{2}{\beta}\cdot\frac{1}{m}\sum_{i=1}^{m}\log\sigma(\beta\cdot\Delta_{i})+\alpha\cdot\frac{1}{n}\sum_{j=1}^{n}f(\theta_{t},z_{j})+\mathcal{P}(\xi_{t},\mu)$
23:  end if
24:  $\theta_{t+1}\leftarrow\theta_{t}-\eta\cdot\nabla L(\theta_{t})$
25:end for
As shown in Algorithm 1, we first compute the average token embedding for each sample in the retain dataset using the pre-unlearning model, thereby obtaining a sentence-level embedding. A global retain embedding is derived by averaging the sentence-level embeddings over all retain samples. Notably, the retain embedding only needs to be computed once initially and can be incrementally updated for multiple unlearning requests. Specifically, as the retain dataset changes with each forgetting request, the global average embedding can be efficiently adjusted by subtracting the embedding of the forget samples from the original average, instead of recomputing all embeddings from scratch. In each iteration of the unlearning process, we compute the average token embedding for each forget sample and estimate its entanglement coefficient by measuring its similarity to the global retain embedding, thereby enabling fine-grained adjustment of the forgetting effort. This coefficient can then reweight the loss across various unlearning optimizers.
Algorithm 2 EAGLE ($E_{\mathcal{D}_{r}}$$\xi_{t}$)
1:Input: Retain embedding $E_{\mathcal{D}_{r}}$, forget samples $\xi_{t}$, entanglement temperature $k$
2:Output: Normalized similarity weights $\{w_{i}\}_{z_{i}\in\xi_{t}}$
3:for $z_{i}\in\xi_{t}$ do
4:  $E_{i}^{f}\leftarrow\text{AvgTokenEmbedding}(\theta_{t},z_{i})$
@@
5:  $cos_{i}\leftarrow\langle E^{f}_{i},E^{r}\rangle/\left(\|E^{f}_{i}\|_{2}\cdot\|E^{r}\|_{2}\right)$
6:  $w_{i}\leftarrow\exp(-cos_{i})$
7:end for
[…]
9:return  $\{w_{i}\}_{z_{i}\in\xi_{t}}$
@@
The specific method for computing the entanglement-based unlearning coefficient is detailed in Algorithm 2. Specifically, we first calculate the cosine similarity between the forget and retain embeddings, which ranges from -1 to 1. To ensure the coefficient remains positive and captures the negative correlation, we apply an exponential transformation $\exp(-cos)$, so that lower similarity (closer to -1) yields a larger coefficient. In this way, forget samples that are less similar to the retain dataset in the embedding space are assigned larger coefficients and thus receive stronger forgetting to better align with the behavior of the retrain model. Intuitively, when a forget sample closely resembles the retain dataset, the retrain model, which is trained without the forget sample, is likely to perform well on it, as the retain knowledge naturally generalizes to similar inputs. In the extreme case where a forget sample is almost identical to part of the retain dataset (e.g., a paraphrased sentence such as “X is born in Y” rewritten as “The birthplace of X is Y”), the retrain model will naturally retain high performance on it. Therefore, samples with lower similarity to the retain data require stronger forgetting to remove residual knowledge and better approximate the behavior of the retrain model. Finally, the coefficients are normalized via softmax operation with a temperature parameter $k$ to flexibly control the forgetting effort and achieve adaptable unlearning behavior. Overall, EAGLE provides a lightweight, entanglement-aware framework that guides loss reweighting, leading to more precise and stable unlearning.
5.2 EAGLE with Proxy Constraint (EAGLE-PC)
While entanglement-awareness guided loss reweighting (EAGLE) adaptively modulates the forgetting effort across samples based on embedding and token-level similarity, gradient-based unlearning methods still suffer from a fundamental limitation: unbounded unlearning. Unlike gradient descent with natural convergence, gradient ascent lacks a stable stopping point. As a result, continued updates can lead to excessive deviation in the model’s predictions on forget samples, causing them to diverge from responses to naturally unseen or semantically similar inputs. This over-forgetting undermines generalization and introduces distributional inconsistencies that increase the risk of unlearning detection.
Setting an appropriate stopping criterion is critical to mitigating over-forgetting during the unlearning process. Since gradient ascent-based unlearning lacks a well-defined point of convergence, it is inherently difficult to determine when to stop. It is non-trivial to identify a suitable threshold in practice since it can vary substantially across model architectures, training datasets, and the degree of entanglement between forget and retain data. A fixed threshold may be too conservative for some samples and too aggressive for others, leading to either under-forgetting or over-forgetting.

### v2

@@ 3.2. Mathematical Modeling for Unlearning
    (2) $L_{\text{GA}}(\theta^{\prime},\mathcal{D}_{f})=-\mathbb{E}_{\mathcal{D}_{f}}[\log(f(\theta^{\prime},y|x))],$
    where the optimization is performed on the forget dataset $\mathcal{D}_{f}$ without referencing the retain dataset $\mathcal{D}_{r}$. Notably, the target label $y$ used in this objective can be flexibly chosen. Instead of the original correct output, one may substitute it with an arbitrary response $\widetilde{y}$ that the model is encouraged to output post-unlearning, such as random answers (Eldan and Russinovich, 2023), or simply answering "I don’t know" (Maini et al., 2024).
@@
    (3) $L(\theta^{\prime},\mathcal{D}_{r})=-\mathbb{E}_{\mathcal{D}_{r}}[\log(f(\theta^{\prime},y|x))],$
@@
  • Negative Preference Optimization (NPO)(Zhang et al., 2024): NPO performs preference optimization using only negative responses as supervision signals. The forget loss is formulated as:
    (4) $L_{\text{NPO},\beta}(\theta^{\prime})=-\frac{2}{\beta}\mathbb{E}_{\mathcal{D}_{f}}\left[\log\sigma(-\beta\log\frac{f(\theta^{\prime},y|x)}{f(\theta^{o},y|x)})\right],$
    where $\sigma(t)=1/(1+e^{-t})$ is the sigmoid function, $\beta>0$ is the inverse temperature and $f(\theta^{o})$ is the model trained on $\mathcal{D}_{\text{train}}$. NPO+GD combines the NPO loss with the retain loss from GD.
3.3. Unlearning Benchmark
In large language model unlearning research, evaluating a model’s ability to forget specific data while preserving overall performance is crucial. To standardize such evaluations, several benchmarks have been proposed (Maini et al., 2024; Li et al., 2024; Shi et al., 2024), among which TOFU (Maini et al., 2024) and MUSE (Shi et al., 2024) are the most representative. TOFU evaluates the model’s ability to forget specific data using around 200 fictitious author profiles, comparing outputs between the retrained and unlearned models to assess both forget quality and model utility. MUSE provides a comprehensive, multi-dimensional evaluation across six properties, including knowledge memory (knowmem), verb memory (verbmem), and privacy leakage (privleak). testing models on large-scale datasets such as books and news articles. These benchmarks together provide a more standardised framework for assessing unlearning methods in large language models, enabling quantitative comparison of unlearning effectiveness and utility preservation.
3.4. Memorization Score
Deep neural networks inherently memorize their training data. Feldman et al. (Feldman, 2020a) show that such instance-level memorization is essential for generalization under long-tailed distributions. This perspective underpins the use of memorization scores below as a proxy to quantify the degree to which individual samples are preserved in the model’s memory.
Assumption 1 (Memorization Score(Feldman, 2020a)).
@@ 3.4. Memorization Score
(5) $\text{mem}(\mathcal{A},\mathcal{D},i)=\Pr_{f\sim\mathcal{A}(\mathcal{D})}\left[f(x_{i})=y_{i}\right]-\Pr_{f\sim\mathcal{A}(\mathcal{D}\setminus{i})}\left[f(x_{i})=y_{i}\right],$
@@
4. Limitations of Existing Unlearning Methods and Our Insights
Neural networks do not memorize all samples equally, some examples are deeply memorized, while others are easily forgotten (Feldman, 2020b; Toneva et al., 2019). Consequently, effective unlearning requires adaptive control over the unlearning process across different dimensions. However, existing approaches often fail to achieve such fine-grained control, leading to inefficiency or over-unlearning. We summarize these challenges from three complementary perspectives: sample-level, iteration-level, and overall-level unlearning effort.
Sample-level: determining how much to forget for each sample. In practice, we first observe that unlearning effort should varies substantially across samples: some examples are tightly entangled with retain data, while others can be safely removed without affecting model utility. As shown in Figure 1, samples with higher entanglement tend to exhibit stronger memorization behaviors, indicating that the model has overfitted to those specific instances, which suggests that entanglement can serve as a useful signal for estimating how much each sample should be forgotten. Specifically, the update induced by EGU exhibits a smaller Euclidean distance (1.0901 vs. 1.2892), suggesting that its update magnitude is also closer to that of retraining. Instead of explicitly measuring memorization, which typically requires retraining or influence-based analysis, we introduce an entanglement-guided loss reweighting mechanism that provides a lightweight and effective alternative. As shown in Figure 2, EGU achieves better unlearning performance while significantly reducing computational overhead, enabling efficient and fine-grained control over sample-level unlearning.
@@ 4. Limitations of Existing Unlearning Methods and Our Insights
Figure 1. Normalized distributions of different unlearning strategies and memorization score. All distributions are min-max normalized for fair comparison. EGU most closely approximates the memorization score distribution.
Figure 2. Comparison of different unlearning methods in terms of performance and computational cost under 5% forgetting rate on LLaMA2-7B with retain dataset.
Figure 3. Cross-batch interference and representation drift during unlearning: as unlearning proceeds, loss on untouched samples rises while embedding similarity declines, indicating unintended over-unlearning.
Iteration-level: adapting unlearning effort over time. As shown in Figure 3, updates applied to one forget batch inevitably propagate to others through the shared parameter space, leading to cross-batch interference. As unlearning progresses, we observe a continuous increase in the loss of untouched samples, together with a steady decline in embedding similarity. This indicates a global representation drift rather than a localized adjustment, meaning that unlearning effects are not confined to targeted samples. These dynamics show that LLM unlearning is inherently coupled across samples and optimization steps, where parameter shifts can propagate and amplify, increasing the risk of over-unlearning and destabilizing the representation space. Although such representation shifts are generally undesirable side effects of unlearning, they provide valuable signals that, if properly exploited, can be used to more effectively guide and regulate the unlearning process.
Figure 4. Forgetting dynamics at 1%, 5%, and 10% comparing GD (dashed) with EGUP_w/o (GD) (solid). Standard GD occasionally leads to excessive unlearning.
Figure 5. Overview of the EGUP framework, which integrates dual entanglement‑guided unlearning with a proxy constraint to steer and stabilize the unlearning process, ensuring effective forgetting while preserving essential knowledge.
Overall-level: determining when to stop unlearning. A key challenge in unlearning lies in the fact that the target degree of unlearning is inherently undefined. During the unlearning process, the model has no clear indication of whether the current level of unlearning is sufficient or already excessive. As shown in Figure 4, the forget loss keeps increasing during the unlearning process and eventually exceeds its corresponding loss under the retrained model, indicating that the model continues to push forget samples away even beyond what is necessary for effective unlearning. This uncontrolled behavior often leads to over-unlearning, where certain samples are excessively altered beyond what is necessary. Therefore, it becomes essential to design mechanisms that enable the model to adaptively recognize when to stop unlearning, achieving a balance between effective removal and distributional stability.
5. Methodology
In real-world scenarios, models do not memorize all samples equally, making it difficult to determine how much unlearning effort each sample requires. EGUP addresses this by introducing an entanglement-guided perspective, where the unlearning effort is adaptively determined by how strongly a forget sample is entangled with retained knowledge. Yet the model still lacks a reliable behavioral reference indicating when unlearning should stop. EGUP therefore further introduces an external behavioral constraint that approximates the desired “knowledge-removed” outputs, providing a soft stopping signal and stabilizing the unlearning process. The overall workflow of EGUP is illustrated in Figure 5.
5.1. Entanglement-Guided Unlearning (EGU)
In this paper, we propose Entanglement-Guided Unlearning, a scalable and principled framework that enables fine-grained control over the unlearning process. EGU leverages dual entanglement signals to guide unlearning: (1) Inter-sample entanglement quantifies the similarity between each forget embedding and the aggregated embedding of the retain dataset. Samples that are highly entangled with retained knowledge i.e., exhibiting higher embedding similarity are likely to be well-predicted by a retrained model, requiring minimal unlearning effort, whereas samples with lower entanglement are forgotten more aggressively. (2) Intra-sample entanglement captures the temporal evolution of each forget embedding across iterations. By continuously tracking how an embedding shifts across iterations even under cross-sample interference induced by updates on other forget samples within the shared parameter space—it adaptively modulates the unlearning process and promotes a smooth, stable trajectory of progressive unlearning.
@@ 5.1. Entanglement-Guided Unlearning (EGU)
Forget samples that are less entangled with retain dataset are typically semantically distant in the embedding space. These samples encode unique information with little overlap with the retain data, warranting stronger unlearning effort. EGU leverages this inter-sample entanglement signal to guide unlearning, approximating it via the cosine similarity between each forget sample and the aggregated embedding of the retain dataset, and adaptively modulating unlearning effort for more precise and effective removal of the target knowledge. In addition, intra-sample entanglement tracks the temporal evolution of each forget embedding, adaptively shaping the unlearning trajectory to ensure smooth and stable unlearning progresses. Together, inter- and intra-sample entanglement enable EGU to perform precise and stable unlearning. The overall framework is detailed in Algorithm 1. Notably, the framework is designed to be modular and flexible, allowing seamless integration into existing unlearning methods, including gradient ascent (Jang et al., 2022), gradient difference (Yao et al., 2024c), negative preference optimization (NPO) (Zhang et al., 2024), and NPO+GD (NPO with gradient descent on the retain dataset).
Algorithm 1 Entanglement-Guided Unlearning with Proxy Constraint
@@
1:Input: Retain dataset $\mathcal{D}_{r}=\left\{z_{j}\right\}_{j=1}^{|\mathcal{D}_{r}|}=\left\{(x_{j},y_{j})\right\}_{j=1}^{|\mathcal{D}_{r}|}$, Forget dataset $\mathcal{D}_{f}=\left\{z_{i}\right\}_{i=1}^{|\mathcal{D}_{f}|}=\left\{(x_{i},y_{i})\right\}_{i=1}^{|\mathcal{D}_{r}|}$, model loss function $f(\theta)$, retain strength $\alpha$, contrastive strength $\mu$
@@
4:for $t=0$ to T1T-1 do
5:  $E_{\mathcal{D}_{r}}^{t}\leftarrow\frac{1}{|\mathcal{D}_{r}|}\sum_{x_{j}\in\mathcal{D}_{r}}\text{AvgTokenEmbedding}(\theta_{t},x_{j})$
6:  Sample $\xi_{t}=\left\{z_{i}\right\}_{i=1}^{m}$ from $\mathcal{D}_{f}$
7:  for $\{x_{i},y_{i}\}\in\xi_{t}$ do
8:   $E_{x_{i}}^{t}\leftarrow\text{AvgTokenEmbedding}(\theta_{t},x_{i})$
@@
9:   $\mathrm{sim}_{i}\leftarrow\langle E_{x_{i}}^{t},E_{\mathcal{D}_{r}}^{t}\rangle/\left(\|E_{x_{i}}^{t}\|_{2}\cdot\|E_{\mathcal{D}_{r}}^{t}\|_{2}\right)$
10:   $\mathrm{sim}_{i}^{\text{prev}}\leftarrow\langle E_{x_{i}}^{t},E_{x_{i}}^{t-1}\rangle/\left(\|E_{x_{i}}^{t}\|_{2}\cdot\|E_{x_{i}}^{t-1}\|_{2}\right)$
11:   $w_{i}\leftarrow\exp(-\mathrm{sim}_{i})$
12:  end for
13:  $w_{i}\leftarrow\text{softmax}\big(k\cdot w_{i}\big)$ for all $i$
14:  $L_{c}(\theta_{t})=-\log\frac{\exp({\mathrm{sim}_{i}})}{\exp({\mathrm{sim}_{i}})+\exp\left(\mathrm{sim}_{i}^{\text{prev}}\right)}$
15:  $L(\theta_{t})=-\frac{1}{|\mathcal{D}_{f}|}\sum_{i=i}^{|\mathcal{D}_{f}|}w_{i}\cdot f(\theta_{t},z_{i})+\alpha\cdot\frac{1}{|\mathcal{D}_{r}|}\sum_{j=1}^{|\mathcal{D}_{r}|}f(\theta_{t},z_{j})+\mu\cdot L_{c}(\theta_{t})+\mathcal{P}(\xi_{t},\mu)$
16:  $\theta_{t+1}\leftarrow\theta_{t}-\eta\cdot\nabla L(\theta_{t})$
17:end for
As shown in Algorithm 1, the proposed EGUP framework performs unlearning through dual entanglement–aware optimization, guided by both inter- and intra-sample entanglement. At every iteration $t$ we first compute the sentence-level (average token) embedding for each sample in the retain dataset and derive a global retain embedding $E_{\mathcal{D}_{r}}^{t}$ by averaging over all retain samples. the model processes a mini-batch $\xi_{t}$ from the forget dataset. For each forget sample $x_{i}$, we compute its sentence-level embedding $E_{x_{i}}^{t}$ under the current parameters $\theta_{t}$, and measure two types of entanglement: (1) the inter-sample entanglement $\mathrm{sim}_{i}=\cos(E_{x_{i}}^{t},E_{\mathcal{D}_{r}}^{t})$, which quantifies the semantic entanglement between forget and retain representations; and (2) the intra-sample entanglement $\mathrm{sim}_{i}^{\text{prev}}=\cos(E_{x_{i}}^{t},E_{x_{i}}^{t-1})$, which captures the parameter-dependent drift of forget embeddings across iterations.
The overall unlearning objective integrates both inter- and intra-sample entanglement–guided optimization. The term $\frac{1}{|\mathcal{D}_{f}|}\sum_{i}w_{i}\cdot f(\theta_{t},x_{i})$ enforces adaptive unlearning guided by inter-sample entanglement, reflecting the semantic correlation between forget and retain samples. Meanwhile, the $\mu\cdot\log\frac{\exp(\mathrm{sim}_{i})}{\exp(\mathrm{sim}_{i})+\exp(\mathrm{sim}_{i}^{\text{prev}})}$ regularizes intra-sample entanglement, adaptively shaping the unlearning trajectory according to the temporal progression of each sample’s representation. The model parameters are subsequently updated via gradient descent: $\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}L(\theta_{t})$. Depending on the specific unlearning strategy, the loss term $f(\theta_{t},x_{i})$ can instantiate various unlearning objectives such as Gradient Ascent (GA) and Negative Preference Optimization (NPO), enabling flexible adaptation of EGUP to different unlearning paradigms.
@@
To further clarify the motivation behind dual entanglement-guided optimization, note that the cosine similarity between forget and retain embeddings ranges from 1-1 to $1$. To ensure the coefficient remains positive and captures the negative correlation, we apply an exponential transformation $\exp(-cos)$, so that lower similarity (closer to -1) yields a larger coefficient. In this way, forget samples that are less similar to the retain dataset in the embedding space are assigned larger coefficients and thus receive stronger unlearning to better align with the behavior of the retrained model. Intuitively, when a forget sample closely resembles the retain dataset, the retrained model, which is trained without the forget sample, is likely to perform well on it, as the retain knowledge naturally generalizes to similar inputs. In the extreme case where a forget sample is almost identical to part of the retain dataset (e.g., a paraphrased sentence such as “X is born in Y” rewritten as “The birthplace of X is Y”), the retrained model will naturally retain high performance on it. Therefore, samples with lower similarity to the retain data require stronger unlearning to remove residual knowledge and better approximate the behavior of the retrained model. Finally, the coefficients are normalized via softmax operation with a temperature parameter $k$ to flexibly control the unlearning effort and achieve adaptable unlearning behavior.
Building upon this semantic weighting mechanism, the second term, $\mathrm{sim}_{i}^{\text{prev}}=\cos(E_{x_{i}}^{t},E_{x_{i}}^{t-1})$, captures intra-sample entanglement, providing a temporal regularization that tracks and guides the evolution of each forget embedding throughout the unlearning process. In the early stages, when $\mathrm{sim}_{i}^{\text{prev}}$ tends to be large due to minimal parameter updates, the contrastive term yields a stronger gradient, accelerating representation divergence and amplifying the unlearning effect. As unlearning proceeds and embeddings progressively decorrelate, $\mathrm{sim}i^{\text{prev}}$ decreases, causing the contrastive penalty to diminish relative to the weighted unlearning loss. This gradual reduction naturally stabilizes optimization, preventing over-unlearning and maintaining representational coherence. By encouraging moderate divergence between $E_{x_{i}}^{t}$ and $E_{x_{i}}^{t-1}$, the temporal term thus promotes a smooth and consistent embedding trajectory that reflects progressive unlearning.
5.2. EGU with Proxy Constraint (EGUP)
While the entanglement-guided unlearning framework adaptively modulates the unlearning effort across samples and iterations based on inter-sample and intra-sample entanglement, it still faces a fundamental challenge in controlling the overall degree of unlearning. In practice, it is often difficult to determine how much unlearning is sufficient. Without a clear criterion, continued updates can excessively alter the model’s predictions on forget samples, pushing them away from responses to naturally unseen inputs. Such unnecessary divergence not only undermines generalization but also introduces distributional inconsistencies, increasing the risk of unlearning detection. This highlights the need for an appropriate stopping criterion to prevent such undesirable deviations during unlearning. However, gradient-ascent–based unlearning offers no natural point of convergence, making it inherently difficult to decide when the level of unlearning becomes sufficient for the intended objective. The optimal threshold can vary substantially across model architectures, training datasets, and the degree of entanglement between forget and retain data. A fixed threshold may be too conservative for some samples and too aggressive for others, leading to either under-unlearning or over-unlearning.